# How to Collaborate: Towards Maximizing the Generalization Performance in Cross-Silo Federated Learning

### v1

@@ Appendix A Proof of Theorem 1
where (b) follows the fact $x+y\leq|x|+|y|,\forall x,y\in\mathbb{R}$. Besides, (c) and (d) apply the definition of several risks directly. Moreover, we obtain the result in (e) following [56].
@@
Next, we provide an upper bound for the training error $\epsilon_{G_{i}}(h)$. According to Theorem 2 of [57], with probability at least $1-\delta$, we have
@@ Appendix B Additional Lemma
$\leq$ $-\eta^{t}\sum_{i\in\mathcal{N}}\mathbb{E}\langle\nabla J_{i}(\mathbf{w}_{i}^{t+1}),\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t+1})\rangle$
$+\frac{L(\eta^{t})^{2}}{2}\sum_{i\in\mathcal{N}}\mathbb{E}\big{[}\big{\|}\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t+1})\big{\|}^{2}\big{]}+\frac{L(\eta^{t})^{2}N}{2}\sigma^{2}.$ (35)
@@
Using the Assumption 1, we have
@@ Appendix C Proof of Theorem 2
We upper bound the inner product in (35) as follows:
@@
$\overset{(\text{b})}{\leq}$ $-\frac{1}{2}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})\|^{2}]-\frac{1}{2}\mathbb{E}[\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
$+\frac{1}{2}\sum_{j\in G_{i}^{t}}a_{j}^{t}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
$=$ $-\frac{1}{2}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})\|^{2}]-\frac{1}{2}\mathbb{E}[\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
$+\frac{1}{2}\sum_{j\in G_{i}^{t}}a_{j}^{t}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})$
$+\nabla J_{i}(\mathbf{w}_{j}^{t})-\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
@@
$+\frac{1}{4\sqrt{\eta^{t}}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]$
$+\frac{\sqrt{\eta^{t}}}{4}\sum_{j\in G_{i}^{t}}a_{j}^{t}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{j}^{t})-\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
@@
$+\frac{1}{4\sqrt{\eta^{t}}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]$
$+\frac{\sqrt{\eta^{t}}}{4}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2},$ (37)
where (a) follows the fact $\langle\mathbf{x},\mathbf{y}\rangle=\frac{1}{2}\|\mathbf{x}\|^{2}+\frac{1}{2}\|\mathbf{y}\|^{2}-\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|^{2}$, (b) adopts the Jensen’s inequality, and (c) applies the inequality $\|a+b\|^{2}\leq\frac{1}{2\sqrt{\eta^{t}}}\|a\|^{2}+\frac{\sqrt{\eta^{t}}}{2}\|b\|^{2}$, and (d) follows Assumption 3. For the third term in (37), we upper bound it as follows:
$\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]$
@@
$\overset{(\text{f})}{\leq}$ $L^{2}t\sum_{s=0}^{t-1}(\eta^{s})^{2}\mathbb{E}[\|\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}(\mathbf{g}_{z}^{s}-\nabla J_{i}(\mathbf{w}_{z}^{s})+\nabla J_{i}(\mathbf{w}_{z}^{s}))$
@@
$\overset{(\text{g})}{\leq}$ $L^{2}t\sum_{s=0}^{t-1}(\eta^{s})^{2}\mathbb{E}[\|\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}(\nabla J_{i}(\mathbf{w}_{z}^{s})-\nabla J_{j}(\mathbf{w}_{z}^{s}))\|^{2}]$
$+2L^{2}t\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2}$
$\overset{(\text{h})}{\leq}$ $L^{2}t\sum_{s=0}^{t-1}(\eta^{s})^{2}\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}\kappa_{i,z}^{2}+2L^{2}t\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2},$ (38)
where (e) follows Assumption 1, and (f) follows the Jensen’s inequality. (g) and (h) hold due to Assumptions 2-3, respectively.
We plug the result of (37) back into the right-hand side (RHS) of (35) as follows:
@@
$+\frac{\sqrt{\eta^{t}}}{4}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}+\frac{L(\eta^{t})^{2}}{2}\sum_{i\in\mathcal{N}}\mathbb{E}[\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
$+\frac{L(\eta^{t})^{2}N}{2}\sigma^{2}+\frac{\sqrt{\eta^{t}}}{4}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]$
@@
$+\frac{(\eta^{t})^{3/2}}{4}\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}$
$+\frac{1}{4\sqrt{\eta^{t}}}L^{2}t\sum_{s=0}^{t-1}(\eta^{s})^{2}\sum_{i\in\mathcal{N}}\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}\kappa_{i,z}^{2}$
$+\frac{L(\eta^{t})^{2}N}{2}\sigma^{2}+\frac{1}{2\sqrt{\eta^{t}}}L^{2}Nt\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2}.$ (39)
Now we rearrange the terms and sum up both sides of (39) over $t=0,1,\dots,T$ to obtain the following result:
@@
$+\sum_{t=0}^{T-1}\frac{(\eta^{t})^{3/2}}{4}\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}$
$+\sum_{t=0}^{T-1}\frac{1}{4\sqrt{\eta^{t}}}L^{2}t\sum_{s=0}^{t-1}(\eta^{s})^{2}\sum_{i\in\mathcal{N}}\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}\kappa_{i,z}^{2}$
$+\sum_{t=0}^{T-1}\frac{L(\eta^{t})^{2}N}{2}\sigma^{2}+\sum_{t=0}^{T-1}\frac{1}{2\sqrt{\eta^{t}}}L^{2}Nt\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2},$ (40)
where we use $-(\frac{\eta^{t}}{2}-\frac{L(\eta^{t})^{2}}{2})\sum_{i\in\mathcal{N}}\mathbb{E}[\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]<0$ and $-\mathbb{E}[J_{i}(\mathbf{w}_{i}^{T})]\leq-\mathbb{E}[J_{i}(\mathbf{w}_{i}^{*})]$.

### v2

@@ Appendix A Proof of Theorem 1
where (b) follows the fact $x+y\leq|x|+|y|,\forall x,y\in\mathbb{R}$. Besides, (c) and (d) apply the definition of several risks directly. Moreover, we obtain the result in (e) following [35], which provides an upper bound for the error of a hypothesis on the target domain.
@@
Next, we provide an upper bound for the training error $\epsilon_{G_{i}}(h)$. According to [36, Theorem 2], with probability at least $1-\delta$, we have
@@ Appendix B Additional Lemma
$\leq$ $-\eta^{t}\sum_{i\in\mathcal{N}}\mathbb{E}\langle\nabla J_{i}(\mathbf{w}_{i}^{t}),\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\rangle$
$+\frac{L(\eta^{t})^{2}}{2}\sum_{i\in\mathcal{N}}\mathbb{E}\big{[}\big{\|}\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\big{\|}^{2}\big{]}+\frac{L(\eta^{t})^{2}N}{2}\sigma^{2}.$ (36)
@@
Using Assumption 1, we have
@@
$=$ $\mathbb{E}\langle\nabla J_{i}(\mathbf{w}_{i}^{t}),-\eta^{t}\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\rangle+\frac{L(\eta^{t})^{2}}{2}\mathbb{E}[\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\mathbf{g}_{j}^{t}\|^{2}]$
$=$ $\mathbb{E}\langle\nabla J_{i}(\mathbf{w}_{i}^{t}),-\eta^{t}\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\rangle+\frac{L(\eta^{t})^{2}}{2}\mathbb{E}[\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
$+\frac{L(\eta^{t})^{2}}{2}\mathbb{E}[\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})-\sum_{j\in G_{i}^{t}}a_{j}^{t}\mathbf{g}_{j}^{t}\|^{2}]$
@@ Appendix C Proof of Theorem 2
Upper bounding the inner product. We upper bound the inner product in (36) as follows:
@@
=$\overset{(\text{b})}{=}$ $-\frac{1}{2}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})\|^{2}]-\frac{1}{2}\mathbb{E}[\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
$+\frac{1}{2}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})+\nabla J_{i}(\mathbf{w}_{j}^{t})-\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
@@
$+\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]$
$+\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{j}^{t})-\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
@@
$+\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]+\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2},$ (38)
where (a) follows the fact $\langle\mathbf{x},\mathbf{y}\rangle=\frac{1}{2}\|\mathbf{x}\|^{2}+\frac{1}{2}\|\mathbf{y}\|^{2}-\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|^{2}$, in (b) we plus and minus $\nabla J_{i}(\mathbf{w}_{j}^{t})$ in the last term, (c) employs the inequality $\frac{1}{2}\|\mathbf{x}+\mathbf{y}\|^{2}\leq\|\mathbf{x}\|^{2}+\|\mathbf{y}\|^{2}$, and (d) follows Assumption 3. It is worth noting that by collaborating with similar clients, the last term $\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}$ can be mitigated due to their similar objectives.
For the third term in the RHS of (38), we upper bound it as follows:
$\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t+1})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]$
@@
$\overset{(\text{f})}{\leq}$ $L^{2}\sum_{s=0}^{t-1}(\eta^{s})^{2}\mathbb{E}[\|\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}(\mathbf{g}_{z}^{s}-\nabla J_{i}(\mathbf{w}_{z}^{s})+\nabla J_{i}(\mathbf{w}_{z}^{s}))$
@@
$\overset{(\text{g})}{\leq}$ $L^{2}\sum_{s=0}^{t-1}(\eta^{s})^{2}\mathbb{E}[\|\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}(\nabla J_{i}(\mathbf{w}_{z}^{s})-\nabla J_{j}(\mathbf{w}_{z}^{s}))\|^{2}]$
$+2L^{2}\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2}$
$\overset{(\text{h})}{\leq}$ $L^{2}\sum_{s=0}^{t-1}(\eta^{s})^{2}\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}\kappa_{i,z}^{2}+2L^{2}\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2},$ (39)
where (e) follows Assumption 1, and (f) follows the independence among different iterations. (g) and (h) hold due to Assumptions 2 and 3, respectively.
We plug the results of (38) and (39) back into the RHS of (36) as follows:
@@
$+\eta^{t}\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}+\frac{L(\eta^{t})^{2}}{2}\sum_{i\in\mathcal{N}}\mathbb{E}[\|\sum_{j\in G_{i}^{t}}a_{j}^{t}\nabla J_{j}(\mathbf{w}_{j}^{t})\|^{2}]$
$+\frac{L(\eta^{t})^{2}N}{2}\sigma^{2}+\eta^{t}\sum_{i\in\mathcal{N}}\mathbb{E}[\|\nabla J_{i}(\mathbf{w}_{i}^{t})-\nabla J_{i}(\mathbf{w}_{j}^{t})\|^{2}]$
@@
$+\eta^{t}\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}$
$+\eta^{t}L^{2}\sum_{s=0}^{t-1}(\eta^{s})^{2}\sum_{i\in\mathcal{N}}\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}\kappa_{i,z}^{2}$
$+\frac{L(\eta^{t})^{2}N}{2}\sigma^{2}+2\eta^{t}L^{2}N\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2}.$ (40)
Completing the proof. Now we rearrange the terms and sum up both sides of (40) over $t=0,1,\dots,T-1$ to obtain the following result:
@@
$+\sum_{t=0}^{T-1}\eta^{t}\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}$
$+\sum_{t=0}^{T-1}\eta^{t}L^{2}\sum_{s=0}^{t-1}(\eta^{s})^{2}\sum_{i\in\mathcal{N}}\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}\kappa_{i,z}^{2}$
$+\sum_{t=0}^{T-1}\frac{L(\eta^{t})^{2}N}{2}\sigma^{2}+\sum_{t=0}^{T-1}2\eta^{t}L^{2}N\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2},$ (41)
where we use $-(\frac{\eta^{t}}{2}-\frac{L(\eta^{t})^{2}}{2})<0$ and $-\mathbb{E}[J_{i}(\mathbf{w}_{i}^{T})]\leq-\mathbb{E}[J_{i}(\mathbf{w}_{i}^{*})]$.
@@
Appendix D Proof of Corollary 1
Recall the inequality (26) as:
$\leq$ $\frac{2}{\xi_{T}}\underbrace{\sum_{i\in\mathcal{N}}\left(\mathbb{E}[J_{i}(\mathbf{w}_{i}^{0})]-\mathbb{E}[J_{i}(\mathbf{w}_{i}^{*})]\right)}_{S_{1}}$ (42)
$+\frac{1}{\xi_{T}}\underbrace{\sum_{t=0}^{T-1}\eta^{t}\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2}}_{S_{2}}$
$+\frac{1}{\xi_{T}}\underbrace{\sum_{t=0}^{T-1}\eta^{t}L^{2}\sum_{s=0}^{t-1}(\eta^{s})^{2}\sum_{i\in\mathcal{N}}\sum_{z\in\mathcal{C}_{i}^{s}}a_{z}^{s}\kappa_{i,z}^{2}}_{S_{3}}$ (43)
$+\frac{1}{\xi_{T}}\underbrace{\sum_{t=0}^{T-1}L(\eta^{t})^{2}N\sigma^{2}+\sum_{t=0}^{T-1}\eta^{t}L^{2}N\sum_{s=0}^{t-1}(\eta^{s})^{2}\sigma^{2}}_{S_{4}}.$ (44)
If the learning rates satisfy $\lim_{T\rightarrow\infty}\xi_{T}=\lim_{T\rightarrow\infty}\sum_{t=0}^{T-1}\eta^{t}=\infty$, we have:
$\lim_{T\rightarrow\infty}\frac{S_{1}}{\xi_{T}}=0,$ (45)
as $S_{1}$ is irrelevant of $t$. Besides, the second term satisfies:
$\lim_{T\rightarrow\infty}\frac{S_{2}}{\xi_{T}}=\mathcal{O}\left(\frac{\sum_{t=0}^{T-1}\eta^{t}(\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2})^{p}}{\sum_{t=0}^{T-1}\eta^{t}}\right)=0,$ (46)
since $\eta^{t}=\mathcal{O}\left((\sum_{i\in\mathcal{N}}\sum_{j\in G_{i}^{t}}a_{j}^{t}\kappa_{i,j}^{2})^{1/p}\right)$ with $p>0$. In addition, we have $\lim_{T\rightarrow\infty}\frac{S_{3}}{\xi_{T}}=0$ and $\lim_{T\rightarrow\infty}\frac{S_{4}}{\xi_{T}}=0$. Thus, the RHS of (26) converges to zero when $T\rightarrow\infty$.